40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 15 22:16:09 2012, Jenne, Update, SUS, All oplev gains turned to 0 
    Reply  Mon Nov 19 15:15:22 2012, Jenne, Update, SUS, All oplev gains turned on 
Message ID: 7551     Entry time: Mon Oct 15 22:16:09 2012     Reply to this: 7724
Author: Jenne 
Type: Update 
Category: SUS 
Subject: All oplev gains turned to 0 

Steve has promised to fix up all of the oplevs, but it hasn't happened yet, so I've turned all of the oplev gains to zero, so that when the optics are restored we don't have to quickly click them off.

Oplev values that were changed to zero:

PRM P=0.15, Y=-0.3

SRM P=-2.0, Y=2.0

BS P=0.2, Y=-0.2

ITMY P=2.1, Y=-2.0

ITMX P=1.0, Y=-0.5

ETMX P=-0.2, Y=-0.2

ETMY P=0.5, Y=0.6

Also, PRCL was changed in the LSC input matrix from REFL33I to AS55I, since there is no REFL beam out of the IFO :(

ELOG V3.1.3-